# Exploring Holomorphic Retracts

### v6

@@ Conclusion
This establishes the linear independence of $p$ and $u$. Now we get back to finishing the proof that $Z$ is linear, to which end, we get back to the equation (6.3). Since $p,u$ are linearly independent, we obtain in-particular, $e^{i2\theta}/2=1$. As such a $\theta$ does not exist, this concludes our proof by contradiction to establish that $Z$ is a non-linear retract of $D$. ∎
@@
The next proposition establishes that this exhausts all possible examples of non-linear (one-dimensional) retracts through the origin in the direction of a boundary point that fails to be $\mathbb{C}$-extremal. Indeed, this proposition is an expanded version of theorem 1.12 whose proof will thereby get finished here. Before we get to the details, it is very helpful to introduce a notation for brevity: if $a=(a_{1},\ldots,a_{k})$ is any point in $\mathbb{C}^{k}$ (for some $k\in\mathbb{N}$) and each of $v_{1},v_{2},\ldots,v_{k}$ are vectors in $\mathbb{C}^{N}$, then we define $a\cdot v=a_{1}v_{1}+a_{2}v_{2}+\ldots+a_{k}v_{k}$; indeed this is the ‘$a$-linear’ combination of the vectors $v_{1},\ldots,v_{k}$.
@@
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose that $\partial D$ is convex at $p\in\partial D$ but not $\mathbb{C}$-extremal at $p$. Let $Z$ be an arbitrary one dimensional retract of $D$ passing through 00 in the direction of $p$; let $\rho:D\to Z$ be a retraction map. Then, $Z$ can be be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\tilde{\psi}(t)\cdot v$ where $\tilde{\psi}:\Delta\to\mathbb{C}^{N-1}$ is an analytic disc with $\tilde{\psi}(0)=0$ and where $v$ stands for a list of basis vectors $(v_{1},\ldots,v_{N})$ of $\ker\left(D\rho(0)\right)$. Also, after factoring out $t$, the resulting map $t\mapsto p+\tilde{\psi}(t)\cdot v$ gives an analytic disc in the boundary $\partial D$; indeed, this analytic disc is the radial projection of $Z$ onto the boundary of our balanced domain $D$. We may also view $Z$ as the graph of the map on ${\rm span}(p)\cap D$ given by $tp\to t\tilde{\psi}(t)\cdot v$ taking values in $\ker\left(D\rho(0)\right)$.
@@
with $v_{1},\ldots,v_{N-1}$ being a basis for the $(N-1)$-dimensional subspace $M:=\ker\left(D\rho(0)\right)$. Note that $\psi(0)=0$ and $\psi^{\prime}(0)=0$ because $Z$ passes through 00 in the direction of $p$ and $p,v_{1},v_{2},...,v_{N-1}$ are linearly independent. Note that we can write $\psi(t)=t\widetilde{\psi}(t)$, for some holomorphic map $\widetilde{\psi}$ with $\widetilde{\psi}(0)=0$.
@@
Since $\partial D$ is convex near $p$ and $\widetilde{\varphi}(0)=p$, there exist a smaller subdisc $\Delta^{\prime}$ of $\Delta$ such that $\partial D$ is convex near $\widetilde{\varphi}(t)$ for all $t\in\Delta^{\prime}$. Observe that that $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}$ is a harmonic function on $\Delta^{\prime}$ and $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}\leq 0$ on $\Delta^{\prime}$. Since $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}(0)=0$, by the maximum principle we obtain $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}\equiv 0$ on $\Delta^{\prime}$ Consider the holomorphic function $F:\Delta\to\mathbb{C}$ defined by
@@
where $\widetilde{\varphi}_{j}$ denotes the $j^{\text{th}}$ component of $\widetilde{\varphi}$. Since $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}$ vanishes on $\Delta^{\prime}$, ${\rm Re}(F)$ vanishes on $\Delta^{\prime}$. By the open mapping theorem applied to $F$, we deduce that $F$ vanishes on $\Delta^{\prime}$. It follows then by the identity principle that $F\equiv 0$ on $\Delta$. In particular, ${\rm Re}F=\Phi\circ\widetilde{\varphi}$ vanishes on $\Delta$. Hence we conclude that $image(\widetilde{\varphi})\subset H_{p,\mathbb{R}}$. This finishes the proof of claim-1.
@@
In this claim, we obtain that $F_{\mathbb{C}}(p,\partial D)+p$ is contained in every complex supporting hyperplane to $\partial D$ at $p$ i.e., every complex supporting hyperplane to the boundary at the point $p$ has got to contain the copy of the face $F_{\mathbb{C}}(p,\partial D)$ translated to $p$.
Continuing with the proof of the proposition 6.4, write $\widetilde{\varphi}(t)=p+\widetilde{\phi}(t)$, where $\widetilde{\phi}:\Delta\to\mathbb{C}^{N}$ is the holomorphic map given by $\widetilde{\phi}(t)=\widetilde{\psi}(t)\cdot v$. Let $L_{p}$ denote the smallest $\mathbb{R}$-linear subspace of $\mathbb{C}^{N}$ containing $image({\widetilde{\phi}})$ ($L_{p}$ is of-course uniquely determined as the intersection of all such $\mathbb{R}$-linear subspaces). We now prove that $L_{p}$ is a complex vector space. Note that $L_{p}$ is the solution of some (finite) system of linear equations with real coefficients, where each linear equation can be written in the following form
@@
where $\tilde{V}$ is a complementary subspace to ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$. If $\pi$ denotes the projection of $\mathbb{C}^{N}$ onto ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$, then note that $\pi(Z)=Z$. Indeed, this is just a re-statement of the fact that $Z$ is contained within ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$ i.e., has no $\tilde{V}$-component. From the fact that $Z=image(\widetilde{\varphi})$ and the expression of $\widetilde{\varphi}(t)$, it is now easy to see that $Z$ is the graph of the holomorphic map on ${\rm span}(p)$ given by $tp\to t\tilde{\psi}\cdot v$; where $\tilde{\psi}(t)\cdot v$ is a holomorphic map from $\Delta$ into the face $F_{\mathbb{C}}(p,\partial D)$ with the property that $p+\tilde{\psi}(t)\cdot v$ is contained within the connected component of
@@
indeed in its interior, as soon as $\tilde{\psi}\not\equiv 0$. Noting that this connected component is contained within the affine linear projection of $D$ into $p+F_{\mathbb{C}}(p,\partial D)$, we are led to the last statement of the proposition; translating this projection back to the origin, gives a bounded balanced pseudoconvex domain which is a projection of $D$ into the lower dimensional complex linear subspace given by the face $F_{\mathbb{C}}(p,\partial D)$. This finishes the proof in all, both of theorem 1.12 as well as the assertions of this proposition 6.4. ∎
@@
When $N=2$, we may rephrase the conclusion above more precisely as follows: every non-linear retract $Z$ passing through the origin in the direction of $p$ can be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\widetilde{\psi}(t)u$, where $u$ belongs to the face $F_{\mathbb{C}}(p,\partial D)$ and $\widetilde{\psi}$ maps $\Delta$ holomorphically onto bounded domain in $\mathbb{C}$ containing the origin, with $\psi(0)=0$ and $\psi^{\prime}(0)=0$. The proof of this is as follows. First observe that from the proof of above proposition, we obtain there exist non-zero $v\in\mathbb{C}^{N}$ such that $p+\widetilde{\psi}(t)v\in\partial D$ for all $t\in\Delta$. Since $Z$ is non-linear, $\widetilde{\psi}$ is not identically zero. By continuity of $\widetilde{\psi}$ and $\widetilde{\psi}(0)=0$, we get $v\in F_{\mathbb{C}}(p,\partial D)$. It follows from the connectedness of image of $\widetilde{\psi}$, that $\widetilde{\psi}$ maps $\Delta$ into $\widetilde{\Delta}$. By the open mapping theorem, $\widetilde{\psi}$ maps $\Delta$ onto an open subset $\widetilde{\psi}(\Delta)$ within ${\rm int}(\widetilde{\Delta})$.
@@
where $B$ is a strictly convex ball such as $l^{p}$ ball with $p\neq 1,\infty$ and $c=1$. In-fact, we can also let $B$ be a non-convex balanced domain of holomorphy such as the $\ell_{q}$-‘ball’ for $10<q<$ 1 (in which case we must choose $c$ appropriately for instance $c=1/2$) to obtain more concrete examples.
Let us now conclude this section with a couple of remarks. We formulate the first one as the following corollary to highlight that: as a consequence of lemma 6.1 and our proof of theorem 1.10 immediately verifies the truth of that theorem in an important special case.
@@
Let $\pi:=D\rho|_{\rho(0)}$ and $L:=T_{p}Z$. Since $\rho$ is a submersion at the origin, $D\rho_{|_{0}}:\mathbb{C}^{N}\longrightarrow T_{p}Z$ is onto. Next, write
$\mathbb{C}^{N}=Ker(D\rho_{|_{0}})\oplus V_{p},$
where $V_{p}$ is some complemented linear subspace of $Ker(D\rho_{|_{0}})$. Now, the submersivity condition means that $D\rho_{|_{0}}$ when restricted to $V_{p}$ maps $V_{p}$ onto $T_{p}Z$. Since ${\rm dim}(V_{p})=\rm dim(T_{p}Z)$, it follows that $D\rho_{|_{0}}:V_{p}\longrightarrow T_{p}Z$ is a bijective linear map i.e., a linear isomorphism. Now, consider the map
$G=(D\rho_{|_{0}})^{-1}\circ\pi\circ\rho:D_{L}\longrightarrow V_{p}.$
Note then that
$DG_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}\circ\pi\circ D\rho_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}\circ\pi(D\rho_{|_{0}}(v)).$
Further, note that $\pi_{|_{T_{p}Z}}$ is just the identity map. So we get $DG_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}(D\rho_{|_{0}}(v))=v$. We then get by Mazet’s Schwarz lemma that $G$ acts as just the identity map on $D_{L}$. Note also that
$(D\rho_{|_{0}})^{-1}\circ\pi\circ\rho\equiv I_{D_{L}}.$
Therefore, $Z={\rm image}(\rho)$ is the graph of holomorphic function over some linear subspace $D_{L}$ of $D$, finishing the proof.
@@
We ought to make a few remarks to settle any first questions that may arise as to whether the result is vacuous(!) i.e., whether there exist retracts not passing through the origin at all. For this let us only recall Lempert’s results which guarantee the existence of one-dimensional retracts $Z$ through any given pair of points $p,q$ which we take to be linearly independent; if $Z$ passes through origin, then the strict convexity hypotheses here leads by Vesentini’s theorem to the conclusion that $Z$ is linear. But then the linear independence of $p,q$ means that ${\rm dim}(Z)>1$; this contradiction finishes the verification about the existence of many retracts not passing through the origin, the matter of the above theorem. This discussion also answers another basic question: are there examples of a bounded balanced domain of holomorphy $D$, wherein there exist pairs of one-dimensional retracts which are not ${\rm Aut}(D)$-equivalent i.e., there is no automorphism carrying one onto the other? In light of the above discussion which actually only requires $\mathbb{C}$-extremality, we may readily cite $D$ to be the $\ell^{1}$-ball in which the origin is fixed by every of its automorphisms; a retract through any pair of linearly independent points in $D$ is therefore not ${\rm Aut}(D)$-equivalent to any of the retracts passing through the origin.
To address other hypotheses of the foregoing theorem briefly, let us remark that a non-trivial holomorphic retraction map is always a submersion except perhaps on a set of measure zero (page $74$ of the book [Hrsch]). Nevertheless, we admit that it would be more significant to ascertain whether the origin falls outside such a measure zero set, for some retraction map $\rho$ as in the above theorem; or, if all retraction maps of such a class of domains is a submersion throughout the domain, making the submersive hypotheses redundant.
Perhaps a more useful remark would be that as a corollary of the proof of this theorem, it will be seen that the strict convexity can be relaxed to $\mathbb{C}$-extremality to obtain the same description of $Z$, as soon as it ascertained that $L\cap\partial D$ has at-least one point which is $\mathbb{R}$-extremal in $\partial D$.
8. Proof of proposition 1.6

### v7

@@ Conclusion
This establishes the linear independence of $p$ and $u$. Now we get back to finishing the proof that $Z$ is non-linear, to which end, we get back to the equation (6.3). Since $p,u$ are linearly independent, we obtain in-particular, $e^{i2\theta}/2=1$. As such a $\theta$ does not exist, this concludes our proof by contradiction to establish that $Z$ is a non-linear retract of $D$.
An alternative proof of this theorem is provided below without the use of complex geodesics.
First, we consider the case that $\partial D$ is convex near $p$. As $\partial D$ is not $\mathbb{C}$-extremal at $p$, there exists a non-zero holomorphic map $\psi:\Delta\to F_{\mathbb{C}}(p,\partial D)$. After rescaling $\Delta$, if needed, we may assume that $\psi(t)$ is a non-vanishing holomorphic map on $\Delta^{*}$. We want to show that there exists $\delta>0$ such that $p+t\psi(t)\in\partial D$ for all $t\in\Delta_{\delta}$, where $\Delta_{\delta}$ denotes the disk centred at the origin of radius $\delta$.
Now, we claim that there exists $\delta>0$ such that, for every $u\in F_{\mathbb{C}}(p,\partial D)\cap S$, we have $p+su\in\partial D$ for all $s\in\Delta_{\delta}$, where $S=\partial\mathbb{B}_{\ell_{1}}$ denote the unit sphere in the $\ell_{1}$ norm. To prove this, let $\text{co}(D)$ denote the convex hull of $D$. Note that $F_{\mathbb{C}}(p,\partial D)=F_{\mathbb{C}}(p,\partial(\text{co}(D)))$ because $D$ is convex near $p$. For any $u\in F_{\mathbb{C}}(p,\partial(\text{co}(D)))\cap S$ can be written as a convex combination of vectors $u_{j}\in F_{\mathbb{C}}(p,\partial(\text{co}(D)))\cap S$. i.e. $u=c_{1}u_{1}+\ldots+c_{k}u_{k}$, where $c_{j}$’s are non-negative reals with $c_{1}+c_{2}+\ldots+c_{k}=1$. Since each $u_{j}\in F_{\mathbb{C}}(p,\partial(\text{co}(D)))$, there exists $\epsilon_{j}>0$ such that $p+su_{j}\in\partial\text{co}(D)$ for all $s\in\Delta_{\epsilon_{j}}$. Let $\delta:=\min\{\epsilon_{1},\ldots,\epsilon_{k}\}$. Then for each $s\in\Delta_{\delta}$,
$p+su=p+s(c_{1}u_{1}+\ldots+c_{k}u_{k})=c_{1}(p+su_{1})+\ldots+c_{k}(p+su_{k})$
Each term $p+su_{j}\in\overline{\text{co}(D)}$, and since $\overline{\text{co}(D)}$ is convex, it follows that $p+su\in\overline{\text{co}(D)}$ for all $s\in\Delta_{\delta}$. Since $D$ is convex near $p$, we may assume that $D$ is convex at $p+su$ for all $s\in\Delta_{\delta}$. By proposition 2.3.1(b) in [Jrncki_invrnt_dst], we have $\gamma_{D}(0;p+su)=h(p+su)$. Applying proposition 2.3.1(d) in [Jrncki_invrnt_dst], we obtain that $h(p+su)=\widehat{h}(p+su)\leq 1$, where $\widehat{h}$ is the Minkowski function of $\text{co}(D)$. This implies that $p+su\in\overline{D}$ for all $s\in\Delta_{\delta}$. From the proof of proposition 2.17, we obtain that $p+su\in\partial D$ for all $s\in\Delta_{\delta}$, which finishes the proof of the claim.
Since $F_{\mathbb{C}}(p,\partial D)$ is closed under scalar multiplication, for each $t\in\Delta^{*}$, we have
$\frac{\psi(t)}{\|\psi(t)\|}\in F_{\mathbb{C}}(p,\partial D)\cap S.$
Since $\phi(t)=t\|\psi(t)\|$ is continuous on $\Delta^{*}$, there exists $\delta_{0}>0$ such that $\|t\psi(t)\|\leq\delta$ for all $t\in\Delta_{\delta_{0}}$. Hence by the above claim, for all $t\in\Delta_{\delta_{0}}^{*}$, we have
$p+t\psi(t)=p+t\|\psi(t)\|\left(\frac{\psi(t)}{\|\psi(t)\|}\right)\in\partial D$
Define the holomorphic map $\tilde{\varphi}:\Delta\to\mathbb{C}^{N}$ by $\tilde{\varphi}(t)=p+\delta t\psi(\delta t)$. Now, we consider the map $\varphi:\Delta\to D$ defined by $\varphi(t)=t\tilde{\varphi}(t)$. Hence $Z:=\text{image}(\varphi)$ is the graph of the holomorphic map $\phi(t)=\delta t^{2}\psi(\delta t)$ over the linear subspace $\operatorname{span}(p)\cap D$ of $D$. By proposition 3.14, it follows that $Z$ is a retract of $D$.
Next, we assume the weaker and general assumption that $\partial D$ is convex at $p$, together with the Minkowski functional of $D$ being continuous near $p$. Choose any holomorphic map $\tilde{\varphi}:\Delta\to(p+F_{\mathbb{C}}(p,\partial D))\cap\partial D$ , since $\partial D$ is not $\mathbb{C}$-extremal at $p$, there exists atleast one such holomorphic map $\tilde{\varphi}$. So, we write $\tilde{\varphi}(t)=p+\tilde{\psi}(t)$, where $\tilde{\psi}\in\mathcal{O}(\Delta)$ with $\tilde{\psi}(0)=0$ and $\tilde{\psi}(t)\in F_{\mathbb{C}}(p,\partial D)$ for all $t\in\Delta$. Now, we consider the holomorphic map $\varphi:\Delta\to\mathbb{C}^{N}$ defined by $\varphi(t)=t\tilde{\varphi}(t)$. In the above proof, we showed that $\varphi$ maps $\Delta$ into $D$. Note that $\text{span}(p)\cap\text{image}(\tilde{\psi})=\{0\}$ because if $u$ is a non-zero vector in $F_{\mathbb{C}}(p,\partial D)$ then $p$ and $u$ are linearly independent (as shown in the above proof; look at the paragraph following equation(6.3)). This implies that $Z=\text{image}(\varphi)$ is a graph of the holomorphic map $\phi(t)=t\tilde{\psi}(t)$ over the linear retract $\text{span}(p)\cap D$ of $D$. By theorem 3.14, it follows that $Z$ is a retract of $D$.
@@
The next proposition establishes that this exhausts all possible examples of non-linear (one-dimensional) retracts through the origin in the direction of a boundary point that fails to be $\mathbb{C}$-extremal. Indeed, this proposition is an important special case of theorem 1.14 whose proof will be based on this and we therefore deal with it first. Moreover, this gives finer information about the one-dimensional retracts than theorem 1.14, which is an assertion about retracts of all dimensions. Before we get to the details, it is very helpful to introduce a notation for brevity: if $a=(a_{1},\ldots,a_{k})$ is any point in $\mathbb{C}^{k}$ (for some $k\in\mathbb{N}$) and each of $v_{1},v_{2},\ldots,v_{k}$ are vectors in $\mathbb{C}^{N}$, then we define $a\cdot v=a_{1}v_{1}+a_{2}v_{2}+\ldots+a_{k}v_{k}$; indeed this is the ‘$a$-linear’ combination of the vectors $v_{1},\ldots,v_{k}$.
@@
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose that $\partial D$ is convex at $p\in\partial D$ but not $\mathbb{C}$-extremal at $p$. Let $Z$ be an arbitrary one-dimensional retract of $D$ passing through 0 in the direction of $p$; let $\rho:D\to Z$ be a retraction map. Then, $Z$ can be be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\tilde{\psi}(t)\cdot v$ where $\tilde{\psi}:\Delta\to\mathbb{C}^{N-1}$ is an analytic disc with $\tilde{\psi}(0)=0$ and where $v$ stands for a list of basis vectors $(v_{1},\ldots,v_{N-1})$ of $\ker\left(D\rho(0)\right)$. Also, after factoring out $t$, the resulting map $t\mapsto p+\tilde{\psi}(t)\cdot v$ gives an analytic disc in the boundary $\partial D$; indeed, this analytic disc is the radial projection of $Z$ onto the boundary of our balanced domain $D$. We may also view $Z$ as the graph of the map on ${\rm span}(p)\cap D$ given by $tp\to t\tilde{\psi}(t)\cdot v$ taking values in $\ker\left(D\rho(0)\right)$.
@@
with $v_{1},\ldots,v_{N-1}$ being a basis for the (N1)(N-1)-dimensional subspace $M:=\ker\left(D\rho(0)\right)$. Note that $\psi(0)=0$ and $\psi^{\prime}(0)=0$ because $Z$ passes through 0 in the direction of $p$ and $p,v_{1},v_{2},...,v_{N-1}$ are linearly independent. Note that we can write $\psi(t)=t\widetilde{\psi}(t)$, for some holomorphic map $\widetilde{\psi}$ with $\widetilde{\psi}(0)=0$.
@@
Since $\partial D$ is convex near $p$ and $\widetilde{\varphi}(0)=p$, there exist a smaller subdisc $\Delta^{\prime}$ of $\Delta$ such that $\partial D$ is convex near $\widetilde{\varphi}(t)$ for all $t\in\Delta^{\prime}$. Observe that that $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}$ is a harmonic function on $\Delta^{\prime}$ and $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}\leq 0$ on $\Delta^{\prime}$. Since $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}(0)=0$, by the maximum principle we obtain $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}\equiv 0$ on $\Delta^{\prime}$ . Consider the holomorphic function $F:\Delta\to\mathbb{C}$ defined by
@@
where $\widetilde{\varphi}_{j}$ denotes the $j^{\text{th}}$ component of $\widetilde{\varphi}$. Since $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}$ vanishes on $\Delta^{\prime}$, ${\rm Re}(F)$ vanishes on $\Delta^{\prime}$. By the open mapping theorem applied to $F$, we deduce that $F$ vanishes on $\Delta^{\prime}$. It follows then by the identity principle that $F\equiv 0$ on $\Delta$. In particular, ${\rm Re}F=\Phi\circ\widetilde{\varphi}$ vanishes on $\Delta$. Hence we conclude that $image(\widetilde{\varphi})\subset H_{p,\mathbb{R}}$. This finishes the proof of claim-1.
@@
As a consequence of this claim, it follows that $F_{\mathbb{C}}(p,\partial D)+p$ is contained in every complex supporting hyperplane to $\partial D$ at $p$ i.e., every complex supporting hyperplane to the boundary at the point $p$ has got to contain the copy of the face $F_{\mathbb{C}}(p,\partial D)$ translated to $p$.
[…]
@@
[…]
@@
indeed in its interior, as soon as $\tilde{\psi}\not\equiv 0$. Noting that this connected component is contained within the affine linear projection of $D$ into $p+F_{\mathbb{C}}(p,\partial D)$, we are led to the last statement of the proposition; translating this projection back to the origin, gives a bounded balanced pseudoconvex domain which is a projection of $D$ into the lower dimensional complex linear subspace given by the face $F_{\mathbb{C}}(p,\partial D)$. This finishes the proof of the proposition (which was for one-dimensional retracts).
Using the foregoing proposition and the arguments in its proof, it is possible to retracts of arbitrary dimension but we do this at the end of this section, after the following remarks and examples.
@@
[…]
@@
where $B$ is a strictly convex ball such as $l^{p}$ ball with $p\neq 1,\infty$ and $c=1$. In-fact, we can also let $B$ be a non-convex balanced domain of holomorphy such as the $\ell_{q}$-‘ball’ for $0<q<1$ (in which case we must choose $c$ appropriately for instance $c=1/2$) to obtain more concrete examples.
Let us now conclude this chapter with a couple of remarks and more importantly, proof of theorem 1.14. We formulate the first one as the following corollary to highlight that: as a consequence of lemma 6.1 and our proof of theorem 1.12 immediately verifies the truth of that theorem in an important special case.
@@
As mentioned above, we shall now use the foregoing proposition 6.4 and the arguments in its proof, to extend it to retracts of arbitrary dimension as stated in theorem 1.14, which we now prove.
Proof of theorem 1.14: Observe that we may apply (the already proven) theorem 1.15 which applies to the setting of the theorem being proved. We can therefore express $Z$ as a graph of a holomorphic map $F$ from $D_{L}$ into $D_{M}$, where $M=\ker(D\rho_{|_{0}})$; notice that $\mathbb{C}^{N}=L\oplus M$. We then pick any point $p$ in $L\cap\partial D$ and consider the retract over the segment spanned by $p$ i.e., let $Z_{p}$ denote the graph of $F$ over $D_{p}:=D\cap{\rm span}(p)$ which is a linear retract precisely by virtue of the convexity hypothesis at $p$. This makes it clear that $Z_{p}$ can be expressed as $\{tp+F(tp):t\in\Delta\}$ and also that $Z_{p}$ is a retract. The foregoing proposition and the arguments in its proof are therefore applicable to this one-dimensional retract $Z_{p}$; we thereby get that $F_{|_{D_{p}}}$ takes values within the complex face $F_{\mathbb{C}}(p,\partial D)$. Viewing $L$ as the union of one-dimensional subspaces, we obtain that $F$ takes values in $\bigcup_{p\in\partial D}F_{\mathbb{C}}(p,\partial D)$. ∎
@@
Consider the holomorphic map $G:D_{L}\to\mathbb{C}^{N}$ defined by the composition of the restricted mappings:
$G(z)=D\rho_{|_{0}}\circ\rho_{|_{D_{L}}}(z).$
Note that $G$ actually maps into $D_{L}$: as $\rho$ maps $D$ into itself and $D\rho_{|_{0}}$ is a linear projection mapping $D$ onto $D_{L}$, it follows that $G$ maps $D_{L}$ into $D_{L}$. Note that for all $z\in D_{L}$, we have
$DG_{|_{0}}(z)=D\rho_{|_{0}}\circ D\rho_{|_{0}}(z)=D\rho_{|_{0}}(z)=z.$
By the hypotheses that $D_{L}$ is convex, we can choose $v\in D_{L}$ such that $v$ is an $\mathbb{R}$-extremal boundary point in $D_{L}$. Note that $DG_{|_{0}}(v)=v$. By Mazet Schwarz lemma 3.5 applied to $G:D_{L}\to D_{L}$, we obtain that $G(z)=z$ for all $z\in D_{L}$. This means that $D\rho_{|_{0}}\circ\rho_{|_{D_{L}}}$ is the identity map on $D_{L}$. Hence we conclude that $Z$ is a graph of a holomorphic function over $D_{L}$, finishing the proof of theorem 1.19.
@@
We ought to make a few remarks to settle any first questions that may arise as to whether the result is vacuous(!) i.e., whether there exist retracts not passing through the origin at all. For this let us only recall Lempert’s results which guarantee the existence of one-dimensional retracts $Z$ through any given pair of points $p,q$ which we take to be linearly independent. Such a $Z$ cannot pass through origin, for otherwise, the strict convexity hypotheses here leads by Vesentini’s theorem to the conclusion that $Z$ is linear. But then the linear independence of $p,q$ means that ${\rm dim}(Z)>1$, a contradiction; this finishes the verification about the existence of many retracts not passing through the origin, the matter of the above theorem. This discussion also answers another basic question: are there examples of a bounded balanced domain of holomorphy $D$, wherein there exist pairs of one-dimensional retracts which are not ${\rm Aut}(D)$-equivalent i.e., there is no automorphism carrying one onto the other? In light of the above discussion which actually only requires $\mathbb{C}$-extremality, we may readily cite $D$ to be the $\ell^{1}$-ball in which the origin is fixed by every of its automorphisms; a retract through any pair of linearly independent points in $D$ is therefore not ${\rm Aut}(D)$-equivalent to any of the retracts passing through the origin. Finally next here, we get to the question of at-least one example wherein all hypotheses of the theorem are satisfied.
Example: Consider the $\ell_{1}$-ball
$B:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<1\}.$
Consider the holomorphic map $\varphi:\Delta\to B$ defined by $\varphi(\zeta)=\left((\frac{1+\zeta}{2})^{2},(\frac{1-\zeta}{2})^{2}\right)$. By proposition 16.2.2 in [Jrncki_invrnt_dst], we obtain that $\varphi$ is a complex geodesic in $B$. Hence
$Z:={\rm image}(\varphi)=\Big{\{}\left((\frac{1+\zeta}{2})^{2},(\frac{1-\zeta}{2})^{2}\right)\;:\;\zeta\in\Delta\Big{\}}$
is a retract of $B$ which we note: does not pass through the origin. Next we want to find a retraction map $\rho$ on $B$ such that $\text{image}(\rho)=Z$. Towards this, it suffices to determine a holomorphic left-inverse for the map $\varphi$ as discussed in proposition 2.32 of the preliminaries section. It is easy to see the map $\phi:B\to\Delta$ defined by $\phi(z,w)=z-w$ serves as the required left-inverse: $\phi\circ\varphi(\zeta)=\zeta$ for all $\zeta\in\Delta$. It follows that the map $\rho(z,w)=\varphi\circ\phi(z,w)$ gives the holomorphic retraction of $B$ onto $Z$. As $B$ is convex with $\mathbb{C}$-extremal boundary, the only hypotheses that needs to be checked is that about the derivative of $\rho$. For this we need its explicit expression which can be computed to be:
$\rho(z,w)=\left(\left(\frac{1+z-w}{2}\right)^{2},\left(\frac{1-z+w}{2}\right)^{2}\right).$
In-particular, $\rho$ maps the origin to $p=(1/4,1/4)$ at which point $L=T_{p}Z$ is given by $L=\{(z,w)\in\mathbb{C}^{2}:z+w=0\}$. Next, we compute the matrix of $D\rho_{|_{(z,w)}}$.
$D\rho_{|_{(z,w)}}=\frac{1}{2}\begin{bmatrix}1+z-w&-(1+z-w)\\ -(1-z+w)&1-z+w\end{bmatrix}$
In particular,
$D\rho_{|_{(0,0)}}=\frac{1}{2}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$
So, $D\rho_{|_{(0,0)}}(z,w)=(\frac{z-w}{2},\frac{-(z-w)}{2})$ which is apparently a projection operator and of norm-one with respect to the $\ell_{1}$-norm. This concludes the verification of all hypotheses of theorem 1.19.
We discuss briefly the form of the conclusion as well. Now we want to show that $Z$ is a graph of a holomorphic function over $B_{L}=B\cap L$, where $L=\{(z,w)\in\mathbb{C}^{2}:z+w=0\}$ which we note is the tangent space to $Z$ at $p=(1/4,1/4)$. Firstly note that $B_{L}=\{(z,-z):|z|<1/2\}$. Consider the holomorphic map $\psi:B_{L}\to B$ defined by
$\psi(z,-z)=\varphi(2z)=\left(\left(\frac{1+2z}{2}\right)^{2},\left(\frac{1-2z}{2}\right)^{2}\right).$
Denote the projection operator $D\rho_{|_{0}}$ by $\pi$ for short. Note that
$\pi\circ\psi_{|_{B_{L}}}(z,-z)=\pi\left(\left(\frac{1+2z}{2}\right)^{2},\left(\frac{1-2z}{2}\right)^{2}\right)=\frac{1}{2}(2z,-2z)=(z,-z)$
The above computation shows that $\pi\circ\psi_{|_{B_{L}}}$ is the identity map on $B_{L}$. This implies that $Z$ is the graph of holomorphic function $\psi$ over $B_{L}$. The reason for the existence of non-linear retractions $\rho$ with $D\rho_{|_{0}}$ being a norm-one projection as in the above example, owes to the boundary failing to be $\mathbb{R}$-extremal (though it is $\mathbb{C}$-extremal). This is brought out by the next result.
Proposition 7.1.
Let $D$ be a bounded balanced strictly convex domain in $\mathbb{C}^{N}$. Let $\rho:D\to D$ be a retraction map on $D$ with $p=\rho(0)$ and $Z:={\rm image}(\rho)$. If $D\rho_{|_{0}}$ is a norm-one projection i.e., a projection which maps $D$ onto $D_{L}$ (where $L=T_{p}Z$ as usual), then $\rho$ is a linear map and hence $Z$ is a linear retract of $D$.
Proof.
Consider the holomorphic map $F:D_{L}\to D$ defined by $F(z)=\rho_{|_{D_{L}}}(z)$. Note that $DF_{|_{0}}(z)=D\rho_{|_{0}}(z)$, which is a linear projection onto $D_{L}$. In particular, $D\rho_{|_{0}}$ is the identity map on $D_{L}$. Applying Mazet’s Schwarz lemma to the map $F:D_{L}\to D$, we obtain that $F(z)=DF_{|_{0}}(z)$. Hence we get that ${\rm image}(F)=D_{L}\subset Z$. Note that $D_{L}$ and $Z$ have the same dimension. Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$. This finishes the proof of the proposition. ∎
8. Proof of theorems 1.51.9 and Corollary 1.11
Before we begin to work out the proof of the pair of theorems as in the title of this section, we need to make some general remarks. Theorems 1.5 is about the irreducibility of balanced $\mathbb{C}$-extremal analytic polyhedra. It will be derived as a consequence of theorem 1.9 about retracts through the origin in such domains; this will therefore be established before theorem 1.5. Along the way, we shall attain useful propositions about biholomorphic inequivalences between co-analytic domains and analytic polyhedra (or more generally, bounded hyperconvex domains). Also included in this section is the proof of corollary 1.11 about the irreducibility of the anisotropic non-convex egg domains, whose boundaries are holomorphically extreme in good contrast to analytic polyhedra; nevertheless, this will be simultaneously attained alongwith the proof of theorem 1.5.